Nuclear power plants often exhaust steam, which you can consider to have the properties of standard air at 5000 ft (since it wants to rise when emitted). Rather than rising, a horizontal wind causes the stack air to move horizontally. What is the horizontal force (in pounds) that the wind imposes on the rising air to cause it to be redirected?6% 3) (23 pts) Nuclear power plants often exhaust steam, which you can consider to have the properties of standard air at 50
Answer #1

Wind force is to be calculated for the given elevation of 5000 feet.

As mentioned in the question a horizontal wind causes the stack air to move horizontally – that is there is a fanning plume formation.

Let there is a stack discharge or stack air emission, Q = 1lb/s.

The velocity of emission vs = is 6ft/s

Force of stack air emission, Fs= Q*vs

Fs= 6 lbf

The resultant force Feq =Fs (in magnitude)

And there is a wind velocity, vw = 15ft/s

Fw = Qw*vw

By using vector algebra of forces

Feq = sqrt (Fw2 + Fs2 - (2* Fw * Fs))

6 = sqrt (Fw2 + 62 - (2* Fw * 6))

36 = Fw2 + 36- 12* Fw

Fw = 12 lbf

Horizontal component of wind force is Fw * cos 450

FwH = 8.485 lbf

media%2F24e%2F24ea54a7-79f8-4cee-afd2-24

Wind force is to be calculated for the given elevation of 5000 feet. As mentioned in the question a horizontal wind causes th